OpenStudy (shreenithaa):

HELP PLEASE a wedding coordinator estimates that Ramirez's wedding will cost approximately $18,000. The wedding coordinator has to stay within $1,500 of the estimated amount. Write the absolute value inequality to show the cost boundaries of the Ramirez's wedding.

OpenStudy (shreenithaa):

@zepdrix

OpenStudy (shreenithaa):

@jennyrlz

zepdrix (zepdrix):

So he has estimated 18000. To be within 1500 of that budget means he can spend `as much as` 18000 + 1500.

zepdrix (zepdrix):

Do you understand how to find the lower value?

OpenStudy (shreenithaa):

guess so..

OpenStudy (shreenithaa):

@zepdrix

OpenStudy (shreenithaa):

how do u so the inequality

zepdrix (zepdrix):

what is the lower amount?

OpenStudy (shreenithaa):

or how do use the words to do it

OpenStudy (shreenithaa):

i guess 1,500

zepdrix (zepdrix):

no

zepdrix (zepdrix):

He can be within 1500 of 18000. So that means he can spend `as much as` 18000 + 1500 and `as little as` _____________?

OpenStudy (shreenithaa):

it doesnt say that maybe its x

zepdrix (zepdrix):

sigh..

OpenStudy (shreenithaa):

is x \[\le\] 18000+1500

zepdrix (zepdrix):

If x is the cost of the wedding, then yes, x has to be less than that amount. That only takes care of half of the inequality though. That's the maximum he can spend. You need to figure out what the minimum amount is before we can write it as an absolute value.

OpenStudy (shreenithaa):

whatever i wrote up there is what my book or hw says

OpenStudy (shreenithaa):

@mathway @mathmate

OpenStudy (shreenithaa):

@jennyrlz

OpenStudy (mathmate):

@shreenithaa Suppose you wanted me to get a bag of chips from the corner store. I asked you how much it would cost. You said between $1.50 and $2.50. ok so far?

OpenStudy (mathmate):

Guess you're not online. I'll continue anyway. This means that the cost is approximately $2.00, plus or minus $0.50. It's because $2.00+0.50 = $2.50, and $2.00-0.50=$1.50. Since you don't know exactly how much, you would give me $2.00 and know that I will either have $0.50 to spare, or I have to add $0.50 to buy the bag of chips. Whether it's more or less than $2.00, you know that the \(difference\) won't exceed $0.50. So you can write If the bag of chip costs x, then |x-$2.00| < $0.50. because the "absolute value" function gives you a positive number whether x is more or less than $2.00. You can apply the same idea to the wedding problem.